OpenStudy (anonymous):

Find the (2n)th term of the G.P. \(-\frac{ 1 }{ 2 }\), 1, -2, 4 ... A. \(2^{2n}\) B. \(-2^{2n}\) C. \(-2^{2n-3}\) D. \(2^{2n-2}\) E. \(-2^{2n-2}\)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large -\frac{1}{2}, 1, -2, 4, ..... \)

OpenStudy (anonymous):

...............?

ganeshie8 (ganeshie8):

find the common ratio r

ganeshie8 (ganeshie8):

r = 4/-2 = -2

OpenStudy (anonymous):

i know

ganeshie8 (ganeshie8):

good:) , use nth term formula

OpenStudy (anonymous):

(-1/2) / [1-(-2)] = -1/6 ??

ganeshie8 (ganeshie8):

\( a_n = a_1 r^{n-1}\)

ganeshie8 (ganeshie8):

n = 2n r = -2

ganeshie8 (ganeshie8):

plugin

OpenStudy (anonymous):

oopsssss

ganeshie8 (ganeshie8):

\(a_{2n} = -\frac{1}{2} (-2)^{2n-1}\)

OpenStudy (anonymous):

(-1/2) * (-2)^(2n-1)

OpenStudy (anonymous):

ok...

ganeshie8 (ganeshie8):

since 2n-1 is odd, that gives - sign

ganeshie8 (ganeshie8):

\(a_{2n} = \frac{1}{2} (2)^{2n-1}\) = \(2^{2n-2}\)

OpenStudy (anonymous):

oic...

ganeshie8 (ganeshie8):

- and - cancelled out

OpenStudy (anonymous):

oh, i know now

ganeshie8 (ganeshie8):

good :)